A car makes a trip due north for three-fourths of the time and due south one-fourth of the time. The average northward velocity has a magnitude of 28 m/s, and the average southward velocity has a magnitude of 24 m/s. Taking northward to be the positive direction, what is the average velocity for the trip?
Answer #1

Displacement in north direction

Y=Vnorth*tnorth =28*(3/4)t=21t

Displacement in south direction

X=Vsouth*tsouth = -24*(1/4)t =-6t

Average velocity

Vavg=Total displacement/Total time =(21t-6t)/t

Vavg=15 m/s